# The $J/\psi\to\phi\eta\pi^{0}$ reaction: $a_{0}(980)$ production, triangle singularity and non $\phi$ background

### v1

@@ Abstract
We study the $J/\psi\to\phi\eta\pi^{0}$ reaction measured recently by the BESIII collaboration with high precision, paying attention to three important aspects: 1) The production of the $a_{0}(980)$ in the $\pi^{0}\eta$ mass distributions, with the typical narrow width observed in isospin violating processes; 2) The origin of two peaks in the $\phi\pi^{0}$ mass distributions that were branded as “non $\phi$” contribution in the experimental analysis; 3) The existence of two triangle mechanisms developing a triangle sigularity at the same energy where the “non $\phi$” contribution peaks in the experiment. However, we also show that the strength of these peaks is very small relative to the observed ones, a feature which is tied to the experimental technique used to identify the $\phi$ looking at $K^{+}K^{-}$ in a narrow window of invariant mass around the $\phi$ mass. We suggest that these triangle singularities could be observed with other methods to identify the $\phi$.
@@ I Introduction
The issue of $f_{0}(980)-a_{0}(980)$ mixing has abstracted much attention [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 4, 24, 25, 26, 27, 28]. The isospin breaking mechanism has been earlier associated to loop effects involving kaons, since both $f_{0}(980)$ and $a_{0}(980)$ couple strongly to $K\bar{K}$, and the mass difference between the charged and neutral kaons induce an isospin violation effect [1, 11]. This picture has become more apparent with the advent of the chiral unitary approach, where the light scalar mesons are dynamically generated from the pseudoscalar-pseudoscalar meson interaction [29, 30, 31, 32]. In particular, both the $f_{0}(980)$ and $a_{0}(980)$ couple strongly to the $K\bar{K}$ component in isospin $I=0$ and $I=1$ respectively.
@@
More work in this direction was made in Ref. [27], where it was shown that in the $J/\psi\to\phi\pi^{0}\eta$ reaction a triangle singularity (TS) should be seen in the $\phi\pi^{0}$ invariant $M_{\phi\pi^{0}}$ around $1385\textrm{ MeV}$, associated to the triangle loop mechanism of Ref. [13] discussed above. Triangle singularities (TSs) appear when in a triangle loop mechanism all intermediate particles can be put on shell and are collinear [36, 37] and in addiction the process can occur at the classical level (Coleman-Norton theorem [38]). A pedagogical and practical approach to TS is given in Ref. [39] and a recent review on this issue can be seen in Ref. [40]. Actually, in isospin violating processes proceeding through triangle mechanisms, the amount of isospin violation is notably enhanced by the presence of a TS [23, 25]. In the present case, since the TS comes from placing intermediate particles on shell, in particular the kaons of the loop, the effects of the mass difference between charged and neutral kaons are magnified with respect to the cases where the particles are off shell, which are more insensitive to the values of the mass.
The idea developed in Ref. [27] stimulated further experimental work, and the BESIII collaboration presented recently results on the $J/\psi\to\phi\pi^{0}\eta$ reaction with high precision in Ref. [28]. In this experiment the $\phi$ was identified by looking at the invariant mass distribution of $K^{+}K^{-}$ pairs, since $\phi\to K^{+}K^{-}$ is the strongest $\phi$ decay mode. Interestingly, a huge peak was observed in the $\phi\pi^{0}$ mass distribution, exactly at the position where a TS was predicted in Ref. [27]. However, the analysis of Ref. [28] identified this peak as a non $\phi$ background contribution, rather than a TS. The idea is that apart from the mechanism leading to $J/\psi\to\pi^{0}\eta\phi,\phi\to K^{+}K^{-}$, which violates isospin, there must be other $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ processes, with $K^{+}K^{-}$ produced in $I=1$, and, hence, without isospin violation, which contribute to the reaction when a cut $(m_{\phi}-10\textrm{ MeV})<M(K^{+}K^{-})<(m_{\phi}+10\textrm{ MeV})$ is performed to identify the $\phi$, as done in Ref. [28]. One might ask whether the appearance of the peak of the non $\phi$ contribution at the same position where the TS is predicted is an accident or a necessary consequence of the dynamics of this non $\phi$ contribution. This is an issue that we want to clarify in the present paper. In addition, we shall also discuss the TS mechanism and link it to the $a_{0}(980)$ production. From this perspective, we shall be able to suggest the type of measurement that should be done to eliminate the non $\phi$ contribution and isolate the TS, which can further clarify the mechanism of isospin violation to produce the $a_{0}(980)$ and the role that the TS plays in it.
@@ II.1 $J/\psi$ decay to a vector and two pseudoscalars
where $\langle\,\rangle$ indicates the trace of these SU(3) matrices. However, arguments of heavy quark symmetry indicate that any extra trace induces a reduction factor of order $\frac{1}{N_{c}}$ [42], where $N_{c}$ is the number of colors. This indicates that the $\langle VPP\rangle$, $\langle V\rangle\langle PP\rangle$ and $\langle VP\rangle\,\langle P\rangle$ structures are the dominant ones and one can neglect the $\langle V\rangle\,\langle P\rangle\langle P\rangle$ structure. In Ref. [27] the $\langle VPP\rangle$ structure is used. In the study of $J/\psi\to\eta h_{1},\pi^{0}b_{1}$ decays in Ref. [43], the $\langle VPP\rangle$ and $\langle V\rangle\langle PP\rangle$ structures were used. Also in the study of $\chi_{c1}\to\eta\pi^{+}\pi^{-}$ in Refs. [44, 45], the term with three traces was found to lead to unacceptable shapes.
@@
where $M_{\rm inv}\equiv M_{\rm inv}(\pi^{0}\eta)$, $G_{K\bar{K}}$ are the loop functions of $K\bar{K}$ in Fig. 1, and $t_{i,j}$ are the scattering matrices of two pseudoscalars evaluated within the chiral unitary approach [29, 30, 31, 32]. We take explicitly into account the $\eta-\eta^{\prime}$ mixing of Ref. [41] in the transition potentials and take the scattering matrices from Ref. [46]. The loop funtions $G_{K\bar{K}}$ are regularized with the cutoff method in Ref. [46] with a cutoff $q_{\rm max}=630\textrm{ MeV}$.
If the masses of $K^{+},K^{-}$ were the same as those of $K^{0},\bar{K}^{0}$, then the two terms in Eq. (II.1) cancel, the $|K\bar{K},I=0\rangle$ state then cannot lead to $\pi^{0}\eta$, and isospin symmetry is preserved. However, if one takes the charged and neutral physical kaon masses, the loops do not cancel leading to a violation of isospin. Actually, since the $t_{i,j}$ matrices are constructed through $T=[1-VG]^{-1}\,V$ and involve the $G$ loops, there is also an isospin violation in the $t_{i,j}$ matrices, and the final isospin violation in $J/\psi\to\phi\pi^{0}\eta$ relies upon the two sources.
@@
which would have a weight $2A_{1}\beta$, with $\beta$ of the order of $N_{c}^{-1}$ according to Ref. [42]. The structure is the same as that of Eq. (4) and hence we automatically can take it into account by substituting in Eq. (5)
@@ II.2 Triangle singularity
Figure 2: Triangle diagrams for $J/\psi\to\phi\pi^{0}\eta$ decay. The momenta of the particles are written in parentheses. $P$ is the sum of the $\pi^{0}$ and $\phi$ momenta. $K^{*}\equiv K^{*}(890)$.
@@
Table 1: $C^{(j)}$ coefficients of Eq. (11). In parenthesis we denote the diagram in Fig. 2.
@@
One can now see, by looking at Eq. (11) and Table 1, that diagrams (a) and (c) of Fig. 2 give equal contribution and diagrams (b) and (d) also provide the same contribution. The two blocks have, however, opposite sign and if the masses of the charged and neutral kaons are equal, and also those for the $K^{*}$, then the four diagrams cancel, as it should be the case for an isospin conserving process. However, once again, if we consider the physical masses of the kaons the cancellation does not occur and we have an isospin violating process.
There is another simplifying method to write the amplitude for the triangle diagram. Since the TS develops when all three intermediate propagator are placed on shell, it is sufficient to take the positive energy part of the propagators. This means that we write
@@
with $\omega(\vec{q}\,)=\sqrt{\vec{q}^{\;2}+m^{2}}$, and take only the first part when $q^{0}$ is positive. With this and taking into account that diagrams (a) and (c) of Fig. 2 are equal and (b) and (d) are also equal and of opposite sign to (a)(c), we can write
@@
where $\omega^{*}(\vec{q}\,)=\sqrt{\vec{q}^{\,2}+m_{K^{*}}^{2}}$, $\omega_{K}(\vec{q}\,)=\sqrt{\vec{q}^{\,2}+m_{K}^{2}}$. Note also that we have considered explicitly the width of the $K^{*}$ by substituting $\omega^{*}(\vec{q}\,)\to\omega^{*}(\vec{q}\,)-i\Gamma_{K^{*}}/2$. The $q^{0}$ integral of Eq. (II.2) is performed using Cauchy integration which picks up $q^{0}=\omega_{K}(\vec{q}\,)$ and then we write, taking into account that
@@
The structure of $t_{c}$ has become very easy, and from there it is easy to see at which $P^{0}$ energy the TS appears, as done in Ref. [39] (see Eq. (18) of that reference). We find that a TS develops at $P^{0}=1385\textrm{ MeV}$, as also found in Ref. [27]. When the $K^{*}$ width is omited, the amplitude develops a singularity, which turns into a pronounced finite peak when the $K^{*}$ width is explicitly considered. We should also note that in all cases $t_{c}$ is logarithmically divergent. However, there is no need to regularize it because we need $t_{c}-t_{n}$ and in the difference the divergence disappears, since for large values of $\vec{q}$ the masses in the integrand do not matter. Next, by taking into account that
@@ II.3 Non $\phi$ contribution in $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ decay
we have unavoidably a contribution from the tree level mechanism associated to Fig. 2, which we depict in Fig. 3.
Figure 3: Tree level contribution for $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ decay (non $\phi$ contribution), with $K^{*}\equiv K^{*}(890)$.
It is clear that the mechanism of Fig. 3 will contribute to the mass distribution when we make the cut of Eq. (21). Note that now the $K^{+}K^{-}$ pair is produced in $I=1$, since there is no isospin violation in that process. Thus, this mechanism does not lead to $\phi$ production but will contribute to the mass spectra of Ref. [28], where indeed a large background below the $\phi$ peak in the $K^{+}K^{-}$ mass distribution is observed. We should also note that since the $J/\psi\to\eta K^{*}\bar{K},\eta\bar{K}^{*}K$ vertices are the same as used to evaluate the TS mechanism, the strength of the mechanisms of Figs. 2 and 3 are related. Yet, since the TS appears from cancellation of diagrams to get the isospin violation mode, which does not occur in the tree level mechanism that does not violate isospin, we can anticipate that the strength of the non $\phi$ contribution should be much larger than that of the TS.
There is another issue worth mentioning here related to Schmid theorem [50] (see also Ref. [51] for detailed discussion concerning this theorem), where the TSs are reabsorbed into the tree level by multiplying it by the phase of the $K^{+}K^{-}$ scattering amplitude in the present case. However, since the tree level and the TS correspond to different isospin states, this should not apply here. In any case the TS contribution is very small here and we would not have the $K^{+}K^{-}$ background if one detected the $\phi$ through a different decay mode, in which case the TS contribution would be the same calculated here.
@@
where the second equation holds after summing over the $K^{*}$ polarizations. Similarly for Fig. 3(b) we get
@@
Then $t_{a_{0}}$ gets converted into $\tilde{t}_{a_{0}}$
@@
and $t_{c}$ gets converted into $\tilde{t}_{c}$
@@ II.4 Tree level with other $K^{*}$ intermediate states
In the BESIII paper it is mentioned that “the components of the non $\phi$ background are very complicated due to $K^{*(l)}$ contributions”, but no further details are provided. Here we have stripped out explicitly one of such contributions in Fig. 3, with the intermediate $K^{*}$ states being the $K^{*}(890)$. In the PDG [52], we find the next two $K^{*}(1^{-})$ resonances, $K^{*}(1410)$ and $K^{*}(1680)$. We shall see that the peak around $2100\textrm{ MeV}$ in the $\phi\pi^{0}$ mass distribution of Ref. [28] is naturally produced with the $K^{*}(1410)$ intermediate state replacing the $K^{*}(892)$ in Fig. 3. The contribution of the $K^{*}(1410)$ is easily done by substituting
@@
    A smaller and wider peak around $2100\textrm{ MeV}$ in the $\phi\pi^{0}$ mass distribution, that we obtain from the tree level amplitude of Fig. 3, with $K^{*}(1410)$ intermediate state;
@@
Figure 4: Results for the $\phi\pi^{0}$ invariant mass distribution. The red error bars represent experimental data from Ref. [28]. The black solid line corresponds to the total result; the green dashed line shows the tree level contribution; the blue dotted line represents the $t_{a_{0}}$ contribution, the brown solid line corresponds to the triangle singularity, and the purple and orange dash-dotted line denote the contributions from the $\phi(1680)$ and $h_{1}(1900)$, repectively. Note that we write $M_{\rm inv}(\phi\pi^{0})$ to keep the same nomenclature as in Ref. [28], but it is actually $M_{\rm inv}(K^{+}K^{-}\pi^{0})$ with the cut of Eq. (21). In the large tree contribution, the $K^{+},K^{-}$ do not correspond to the $\phi$ since they are in $I=1$. This corresponds to the “non $\phi$” contribution of Ref. [28].
In Fig. 4 we show the $\phi\pi^{0}$ mass distribution including all the mechanisms discussed in the work. The data show two clear peaks, one around 1400 MeV and the other around 2100 MeV. The contribution of the two peaks of the $\phi\eta$ mass distribution (see Fig. 7) only affect the $\phi\pi^{0}$ mass distribution at high energy. This allows us to see clearly what is the origin of the two peaks in the $\phi\pi^{0}$ mass distribution. Even if the first peak appears close to the predicted position of a triangle singularity coming from the triangle mechanism considered here (Fig. 2), its origin is different. In fact it comes from the tree level $K^{+}K^{-}$ production of Fig. 3. The peak is reasonably reproduced using a parameter $A_{tT,890}=3.12\times 10^{4}$ to adjust to the counts in the experiment. We also obtain reasonably the second peak by using the same tree level mechanism of Fig. 3 and the $K^{*}(1410)$ resonance, using $A_{t,1410}=6.25\times 10^{4}$ to adjust to the counts in the experiment. This exercise clarifies the origin of these two peaks, that in the experimental analysis of Ref. [28] are attributed to the “non-$\phi$” contribution without further detail. Given the fact that the triangle mechanism of Fig. 2 uses the same input as the tree level with the $K^{*}(890)$ resonance, we can now evaluate its contribution to the $\phi\pi^{0}$ mass distribution in the same scale as Fig. 4. The results are shown in Fig. 5, where we see that, indeed, there is a peak at $1390\textrm{ MeV}$ corresponding to the predicted triangle singularity, but its strength is about $40$ times smaller than the strength of the experimental peak appearing at the same energy. There is a second peak associated to the triangle diagram of Fig. 2, where we consider the intermediate $K^{*}$ as the $K^{*}(1410)$. Its strength is smaller than the one around 1400 MeV and does not show up in Fig. 4. We can see that there is indeed a triangle singularity associated to the second peak of the mass distribution, appearing at the same position as the experimental peak but it is wider and with smaller strength.

### v2

@@ Abstract
We study the $J/\psi\to\phi\eta\pi^{0}$ reaction measured recently by the BESIII Collaboration with high precision, paying attention to three important aspects: 1) the production of the $a_{0}(980)$ in the $\pi^{0}\eta$ mass distributions, with the typical narrow width observed in isospin violating processes; 2) the origin of two peaks in the $\phi\pi^{0}$ mass distributions that were branded as “non-$\phi$” contribution in the experimental analysis; 3) the existence of two triangle mechanisms developing a triangle singularity at the same energy where the “non-$\phi$” contribution peaks in the experiment. However, we also show that the strength of these peaks is very small relative to the observed ones, a feature that is tied to the experimental technique used to identify the $\phi$ looking at $K^{+}K^{-}$ in a narrow window of invariant mass around the $\phi$ mass. We suggest that these triangle singularities could be observed with other methods to identify the $\phi$.
@@ I Introduction
The issue of $f_{0}(980)$-$a_{0}(980)$ mixing has attracted much attention [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 4, 24, 25, 26, 27, 28]. The isospin breaking mechanism has been earlier associated with loop effects involving kaons, since both $f_{0}(980)$ and $a_{0}(980)$ couple strongly to $K\bar{K}$, and the mass difference between the charged and neutral kaons induce an isospin violation effect [1, 11]. This picture has become more apparent with the advent of the chiral unitary approach, where the light scalar mesons are dynamically generated from the pseudoscalar-pseudoscalar meson interaction [29, 30, 31, 32]. In particular, both the $f_{0}(980)$ and $a_{0}(980)$ couple strongly to the $K\bar{K}$ component in isospin $I=0$ and $I=1$ respectively.
@@
More work in this direction was made in Ref. [27], where it was shown that in the $J/\psi\to\phi\pi^{0}\eta$ reaction a triangle singularity (TS) should be seen in the $\phi\pi^{0}$ invariant $M_{\phi\pi^{0}}$ around $1385\textrm{ MeV}$, associated with the triangle loop mechanism of Ref. [13] discussed above. Triangle singularities (TSs) appear when in a triangle loop mechanism all intermediate particles can be put on shell and are collinear [36, 37] and, in addition, the process can occur at the classical level (Coleman-Norton theorem [38]). A pedagogical and practical approach to TS is given in Ref. [39] and a recent review on this issue can be seen in Ref. [40]. Actually, in isospin violating processes proceeding through triangle mechanisms, the amount of isospin violation is notably enhanced by the presence of a TS [23, 25]. In the present case, since the TS comes from placing intermediate particles on shell, in particular, the kaons of the loop, the effects of the mass difference between charged and neutral kaons are magnified with respect to the cases where the particles are off shell, which are more insensitive to the values of the mass.
The idea developed in Ref. [27] stimulated further experimental work, and the BESIII Collaboration recently presented results on the $J/\psi\to\phi\pi^{0}\eta$ reaction with high precision in Ref. [28]. In this experiment, the $\phi$ was identified by looking at the invariant mass distribution of $K^{+}K^{-}$ pairs, since $\phi\to K^{+}K^{-}$ is the strongest $\phi$ decay mode. Interestingly, a huge peak was observed in the $\phi\pi^{0}$ mass distribution, exactly at the position where a TS was predicted in Ref. [27]. However, the analysis of Ref. [28] identified this peak as a non-$\phi$ background contribution, rather than a TS. The idea is that apart from the mechanism leading to $J/\psi\to\pi^{0}\eta\phi,\phi\to K^{+}K^{-}$, which violates isospin, there must be other $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ processes, with $K^{+}K^{-}$ produced in $I=1$, and, hence, without isospin violation, which contribute to the reaction when a cut $(m_{\phi}-10\textrm{ MeV})<M(K^{+}K^{-})<(m_{\phi}+10\textrm{ MeV})$ is performed to identify the $\phi$, as done in Ref. [28]. One might ask whether the appearance of the peak of the non-$\phi$ contribution at the same position where the TS is predicted is an accident or a necessary consequence of the dynamics of this non-$\phi$ contribution. This is an issue that we want to clarify in the present paper. In addition, we shall also discuss the TS mechanism and link it to the $a_{0}(980)$ production. From this perspective, we shall be able to suggest the type of measurement that should be done to eliminate the non-$\phi$ contribution and isolate the TS, which can further clarify the mechanism of isospin violation to produce the $a_{0}(980)$ and the role that the TS plays in it.
@@ II.1 $J/\psi$ decay to a vector and two pseudoscalars
where $\langle\,\rangle$ indicates the trace of these SU(3) matrices. However, arguments of heavy quark symmetry indicate that any extra trace induces a reduction factor of order $\frac{1}{N_{c}}$ [42], where $N_{c}$ is the number of colors. This indicates that the $\langle VPP\rangle$, $\langle V\rangle\langle PP\rangle$, and $\langle VP\rangle\,\langle P\rangle$ structures are the dominant ones and one can neglect the $\langle V\rangle\,\langle P\rangle\langle P\rangle$ structure. In Ref. [27] the $\langle VPP\rangle$ structure is used. In the study of $J/\psi\to\eta h_{1},\pi^{0}b_{1}$ decays in Ref. [43], the $\langle VPP\rangle$ and $\langle V\rangle\langle PP\rangle$ structures were used. Also in the study of $\chi_{c1}\to\eta\pi^{+}\pi^{-}$ in Refs. [44, 45], the term with three traces was found to lead to unacceptable shapes.
@@
where $M_{\rm inv}\equiv M_{\rm inv}(\pi^{0}\eta)$, $G_{K\bar{K}}$ are the loop functions of $K\bar{K}$ in Fig. 1, and $t_{i,j}$ are the scattering matrices of two pseudoscalars evaluated within the chiral unitary approach [29, 30, 31, 32]. We take explicitly into account the $\eta-\eta^{\prime}$ mixing of Ref. [41] in the transition potentials and take the scattering matrices from Ref. [46]. The loop functions $G_{K\bar{K}}$ are regularized with the cutoff method in Ref. [46] with a cutoff $q_{\rm max}=630\textrm{ MeV}$.
If the masses of $K^{+},K^{-}$ were the same as those of $K^{0},\bar{K}^{0}$, then the two terms in Eq. (II.1) cancel, the $|K\bar{K},I=0\rangle$ state then cannot lead to $\pi^{0}\eta$, and isospin symmetry is preserved. However, if one takes the charged and neutral physical kaon masses, the loops do not cancel, leading to a violation of isospin. Actually, since the $t_{i,j}$ matrices are constructed through $T=[1-VG]^{-1}\,V$ and involve the $G$ loops, there is also an isospin violation in the $t_{i,j}$ matrices, and the final isospin violation in $J/\psi\to\phi\pi^{0}\eta$ relies upon the two sources.
@@
which would have a weight $2A_{1}\beta$, with $\beta$ on the order of $N_{c}^{-1}$ according to Ref. [42]. The structure is the same as that of Eq. (4); hence we automatically can take it into account by substituting in Eq. (5)
@@ II.2 Triangle singularity
Figure 2: Triangle diagrams for $J/\psi\to\phi\pi^{0}\eta$ decay. (a), (b), (c), and (d) correspond to different charges in the intermediate states. The momenta of the particles are written in parentheses. $P$ is the sum of the $\pi^{0}$ and $\phi$ momenta. $K^{*}\equiv K^{*}(890)$.
@@
Table 1: $C^{(j)}$ coefficients of Eq. (11). In parentheses we denote the diagram in Fig. 2.
@@
One can now see, by looking at Eq. (11) and Table 1, that Fig. 2(a) and Fig. 2(c) give equal contribution and Fig. 2(b) and Fig. 2(d) also provide the same contribution. The two blocks have, however, opposite sign and if the masses of the charged and neutral kaons are equal, and also those for the $K^{*}$, then the four diagrams cancel, as should be the case for an isospin conserving process. However, once again, if we consider the physical masses of the kaons the cancellation does not occur and we have an isospin violating process.
There is another simplifying method to write the amplitude for the triangle diagram. Since the TS develops when all three intermediate propagators are placed on shell, it is sufficient to take the positive energy part of the propagators. This means that we write
@@
with $\omega(\vec{q}\,)=\sqrt{\vec{q}^{\;2}+m^{2}}$, and take only the first part when $q^{0}$ is positive. With this and taking into account that Fig. 2(a) and Fig. 2(c) are equal and Fig. 2(b) and Fig. 2(d) are also equal and of opposite sign to Fig. 2(a) and Fig. 2(c), we can write
@@
where $\omega^{*}(\vec{q}\,)=\sqrt{\vec{q}^{\,2}+m_{K^{*}}^{2}}$ and $\omega_{K}(\vec{q}\,)=\sqrt{\vec{q}^{\,2}+m_{K}^{2}}$. Note also that we have considered explicitly the width of the $K^{*}$ by substituting $\omega^{*}(\vec{q}\,)\to\omega^{*}(\vec{q}\,)-i\Gamma_{K^{*}}/2$. The $q^{0}$ integral of Eq. (II.2) is performed using Cauchy integration which picks up $q^{0}=\omega_{K}(\vec{q}\,)$ and then we write, taking into account that
@@
The structure of $t_{c}$ has become very easy, and from there it is easy to see at which $P^{0}$ energy the TS appears, as done in Ref. [39] [see Eq. (18) of that reference]. We find that a TS develops at $P^{0}=1385\textrm{ MeV}$, as also found in Ref. [27]. When the $K^{*}$ width is omitted, the amplitude develops a singularity, which turns into a pronounced finite peak when the $K^{*}$ width is explicitly considered. We should also note that in all cases $t_{c}$ is logarithmically divergent. However, there is no need to regularize it because we need $t_{c}-t_{n}$ and in the difference the divergence disappears, since for large values of $\vec{q}$ the masses in the integrand do not matter. Next, by taking into account that
@@ II.3 Non-$\phi$ contribution in $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ decay
we have unavoidably a contribution from the tree level mechanism associated with Fig. 2, which we depict in Fig. 3.
Figure 3: Tree level contribution for $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ decay (non-$\phi$ contribution), with $K^{*}\equiv K^{*}(890)$. (a) and (b) correspond to different charges in the intermediate states.
It is clear that the mechanism of Fig. 3 will contribute to the mass distribution when we make the cut of Eq. (21). Note that now the $K^{+}K^{-}$ pair is produced in $I=1$, since there is no isospin violation in that process. Thus, this mechanism does not lead to $\phi$ production but will contribute to the mass spectra of Ref. [28], where indeed a large background below the $\phi$ peak in the $K^{+}K^{-}$ mass distribution is observed. We should also note that since the $J/\psi\to\eta K^{*}\bar{K},\eta\bar{K}^{*}K$ vertices are the same as used to evaluate the TS mechanism, the strength of the mechanisms of Figs. 2 and 3 are related. Yet, since the TS appears from cancellation of diagrams to get the isospin violation mode, which does not occur in the tree level mechanism that does not violate isospin, we can anticipate that the strength of the non-$\phi$ contribution should be much larger than that of the TS.
There is another issue worth mentioning here related to the Schmid theorem [50] (see also Ref. [51] for detailed discussion concerning this theorem), where the TSs are reabsorbed into the tree level by multiplying it by the phase of the $K^{+}K^{-}$ scattering amplitude in the present case. However, since the tree level and the TS correspond to different isospin states, this should not apply here. In any case, the TS contribution is very small here and we would not have the $K^{+}K^{-}$ background if one detected the $\phi$ through a different decay mode, in which case the TS contribution would be the same calculated here.
@@
where the second equation holds after summing over the $K^{*}$ polarizations. Similarly, for Fig. 3(b) we get
@@
Then $t_{a_{0}}$ gets converted into $\tilde{t}_{a_{0}}$,
@@
and $t_{c}$ gets converted into $\tilde{t}_{c}$,
@@ II.4 Tree level with other $K^{*}$ intermediate states
In the BESIII paper, it is mentioned that “the components of the non-$\phi$ background are very complicated due to $K^{*(l)}$ contributions”, but no further details are provided. Here we have stripped out explicitly one of such contributions in Fig. 3, with the intermediate $K^{*}$ states being the $K^{*}(890)$. From the Particle Data Group [52], we find the next two $K^{*}(1^{-})$ resonances, $K^{*}(1410)$ and $K^{*}(1680)$. We shall see that the peak around $2100\textrm{ MeV}$ in the $\phi\pi^{0}$ mass distribution of Ref. [28] is naturally produced with the $K^{*}(1410)$ intermediate state replacing the $K^{*}(892)$ in Fig. 3. The contribution of the $K^{*}(1410)$ is easily done by substituting
@@
    A smaller and wider peak around $2100\textrm{ MeV}$ in the $\phi\pi^{0}$ mass distribution, which we obtain from the tree level amplitude of Fig. 3, with $K^{*}(1410)$ intermediate state;
@@
Figure 4: Results for the $\phi\pi^{0}$ invariant mass distribution. The red error bars represent experimental data from Ref. [28]. The black solid line corresponds to the total result; the green dashed line shows the tree level contribution; the blue dotted line represents the $t_{a_{0}}$ contribution, the brown solid line corresponds to the triangle singularity, and the purple and orange dash-dotted line denote the contributions from the $\phi(1680)$ and $h_{1}(1900)$, respectively. Note that we write $M_{\rm inv}(\phi\pi^{0})$ to keep the same nomenclature as in Ref. [28], but it is actually $M_{\rm inv}(K^{+}K^{-}\pi^{0})$ with the cut of Eq. (21). In the large tree contribution, the $K^{+},K^{-}$ do not correspond to the $\phi$ since they are in $I=1$. This corresponds to the “non-$\phi$” contribution of Ref. [28].
In Fig. 4 we show the $\phi\pi^{0}$ mass distribution including all the mechanisms discussed in the work. The data show two clear peaks, one around 1400 MeV and the other around 2100 MeV. The contribution of the two peaks of the $\phi\eta$ mass distribution (see Fig. 8) only affect the $\phi\pi^{0}$ mass distribution at high energy. This allows us to see clearly what is the origin of the two peaks in the $\phi\pi^{0}$ mass distribution. Even if the first peak appears close to the predicted position of a triangle singularity coming from the triangle mechanism considered here (Fig. 2), its origin is different. In fact it comes from the tree level $K^{+}K^{-}$ production of Fig. 3. The peak is reasonably reproduced using a parameter $A_{tT,890}=3.12\times 10^{4}$ to adjust to the counts in the experiment. We also obtain reasonably the second peak by using the same tree level mechanism of Fig. 3 and the $K^{*}(1410)$ resonance, using $A_{t,1410}=6.25\times 10^{4}$ to adjust to the counts in the experiment. This exercise clarifies the origin of these two peaks, that in the experimental analysis of Ref. [28] are attributed to the “non-$\phi$” contribution without further detail. Given the fact that the triangle mechanism of Fig. 2 uses the same input as the tree level with the $K^{*}(890)$ resonance, we can now evaluate its contribution to the $\phi\pi^{0}$ mass distribution in the same scale as Fig. 4. The results are shown in Fig. 5, where we see that, indeed, there is a peak at $1390\textrm{ MeV}$ corresponding to the predicted triangle singularity, but its strength is about $40$ times smaller than the strength of the experimental peak appearing at the same energy. There is a second peak associated with the triangle diagram of Fig. 2, where we consider the intermediate $K^{*}$ as the $K^{*}(1410)$. Its strength is smaller than the one around 1400 MeV and does not show up in Fig. 4. We can see that there is indeed a triangle singularity associated with the second peak of the mass distribution, appearing at the same position as the experimental peak but it is wider and with smaller strength.
@@
In Fig. 6, the TS curve is basically invisible. To show its line shape, we plot it in Fig. 7.
Figure 7: Only triangle singularity result for the $\eta\pi^{0}$ invariant mass distribution.
@@
Figure 9: Only triangle singularity result for the $\phi\eta$ invariant mass distribution.